OpenStudy (anonymous):

Which of the following equations represents the greatest value of x? Equation 1 Equation 2 Equation 3 Equation 4 x2 = 24 x3 = 64 x2 = 28 x3 = 125 (5 points) Equation 2 Equation 1 Equation 4 Equation 3

OpenStudy (haleyelizabeth2017):

is that cubed and squared?

OpenStudy (anonymous):

x^2=24 x^3=64 x^2=28 x^3=125

OpenStudy (anonymous):

@chucky00

OpenStudy (haleyelizabeth2017):

okay just making sure.....so we need to sqrt and cuberoot all of these to begin with

OpenStudy (anonymous):

okay

OpenStudy (haleyelizabeth2017):

\[\sqrt 24\] \[\sqrt[3]{64}\] \[\sqrt 28\] \[\sqrt[3]{125}\]

OpenStudy (haleyelizabeth2017):

just do that in the calculator and its simple! :D

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

i think lol

OpenStudy (haleyelizabeth2017):

nope! here is what I got.... 4.9 4 5.3 5 So we can tell it is \(\sqrt 28\)

OpenStudy (anonymous):

oh duh lol!

OpenStudy (anonymous):

so its C!

OpenStudy (haleyelizabeth2017):

lol good job though! Yep

OpenStudy (anonymous):

cool! thanks!

OpenStudy (haleyelizabeth2017):

no problem!